# Y-AR: A Mixed Reality CAD Tool for 3D Wire Bending

### v1

@@ Abstract.
Wire bending is a technique used in manufacturing to mass-produce . CWire bending machines like the DIWire by Pensalabs111https://www.pensalabs.com/diwire-pro have made this process accessible for personal fabrication. However, such machines are controlled using Computer Aided Manufacturing (CAM) software which is hard to use, making custom design challenging. We present Y-AR, a Computer Aided Design (CAD) interface for 3D wire bending. Y-AR uses mixed reality to let designers create structures that physically connect to objects in the environment (such as the mount shown in Figure 1c). The interface incorporates springs as design primitives which (1) apply forces to hold objects, and (2) counter-act dimensional inaccuracies inherently caused by mid air modeling and measurement errors in AR. We demonstrate design workflows to design and fabricate a range of mechanisms designed in Y-AR as well as structures made using free-hand design tools. In our usability evaluation, all 12 participants successfully designed and fabricated a functional bottle holder with Y-AR. 10 out of 12 participants felt that the system aided their design process, rating it above 7 out of 10.
@@ 1. Introduction
Wire bending is a common industrial mass-manufacturing process used to create springs, clips, mounts, and braces. In sculpting, it is used to provide internal support for incomplete structures, particularly in the form of armatures. Most wire bending occurs in mass-production with high-volume, low-mix production. Today, there is a growing trend towards custom wire bending, where each object is heterogeneous and crafted individually rather than the mass manufacture of wire objects with uniform geometry.This custom process is predominantly manual. In the medical sector, it is used to create retainers and braces (Alassiry, 2019), and in puppetry, it enables the production of actuated characters (Xu et al., 2018), both are examples of low-volume, high-mix production (e.g., few unique products).
@@
Machines like the DIWire by Pensalabs have made this process accessible for personal fabrication (Baudisch and Mueller, 2017). Hobbyists have embraced the technology to make connectors, holders, signs, marble tracks, and sculptures. While software exists to operate the machines (e.g., VGPNEXT222https://www.blmgroup.com/en-us/software/vgpnext, Bend-Tech333https://www.bend-tech.com/, and BendPro SIM444https://www.currenttech.com/tube-bending-software/bendpro-sim/), they naturally inherit design approaches from industrial manufacturing. Users apply sequential combinations of bend, rotation, and extrusion operations. Such low-level, machine-specific commands (common in Computer Aided Manufacturing (CAM)) make it hard for users without technical expertise to design structures for wire bending.
Researchers have built software converting existing 3D models to bend paths (Huang et al., 2023; Hsiao et al., 2018). These tools let hobbyists create intricate wire structures based off 3D models, however modifying models after conversion remains challenging, requiring round-trips of conversion. These design tools are agnostic to the strengths and weaknesses of wire bending. Often there are multiple ways to convert 3D models to paths for wire bending, WireRoom (Yang et al., 2021) eloquently addresses this through a user interface that offers multiple generated wire bending paths to choose from. However, to date, there are no Computer Aided Design (CAD) tools designed to support wire bending.
@@
We present Y-AR, a CAD tool for 3D wire bending in mixed reality. It is a fabrication-aware-design (Bermano et al., 2017) tool that lets designers create 3D structures optimized for CNC wire bending. Y-AR overlays designed 3D structures on users’ physical environments in mixed reality to create objects fitting within the physical environment of the designer (examples shown in Figure 2). Y-AR uses design primitives such as mounts, and holders, to connect structures to physical artifacts. Y-AR leverages the elasticity of wires to create mechanisms that apply force to hold objects in place. This springiness compensates for the dimensional inaccuracies inherent to XR-based modeling and hand-tracking. It automatically generates machine instructions for wire benders and assembly instructions for users. We present the design space of primitives consisting of clamps and mounts to hold objects with shapes such as cylinders, boxes, spheres, and higher-order polyhedra. We furthermore demonstrate free-form design (where users draw structures in mid air) by modeling a custom marble track. In our usability evaluation, all 12 participants successfully designed and fabricated a functional bottle holder with Y-AR. 10 out of 12 participants felt that Y-AR aided their design process, rating it above 7 out of 10.
@@ 2. RELATED WORK
This work builds on research on wire bending, fabrication-aware design, and computer-aided design in mixed reality.
@@ 2.1. Wire Bending
Research on wire bending has focused on aesthetic and functional wire bending. For aesthetic approaches, WrapIt created jigs to speed up manual wire bending for jewelry based on 2D patterns (Iarussi et al., 2015). Fabricable 3D Wire Art similarly generate jigs however they base it on 3D models, and images enabling more complex designs (Tojo et al., 2024). With the advent of CNC wire bending machines, there has been an increased focus on advanced wire bending by converting 2D line drawings (Hsiao et al., 2018) or 3D models (Huang et al., 2023) to machine instructions for wire bending. Yeh et al. (2022) propose an algorithm to create 3D wire art with the minimum number of continuous wires. In terms of user interfaces to generate 3D wire art, WireRoom provides an interface to offer multiple options to generate wire bending paths (Yang et al., 2021) allowing users to select their preferred result. PM4WireArt offers users more freedom through a parametric modeling interface designed for conceptual wire art design (Li, 2024). However, instead of focusing on traditional wire bending—where a single wire is bent and rotated into shapes—PM4WireArt treats wire as rigid rods assembled into structures, resembling welded frameworks. This approach facilitates complex design creation highlighting a different use of wire in structural art.
Several projects explore functional instead of purely aesthetic properties of wire bending. For instance, Bend-forming extends the application of wire bending to aerospace engineering. Their work involves the generation of truss structures and explores the concept of deploying wire benders in space to construct large-scale, static structures (Bhundiya and Cordero, 2023). More down to earth, WireFab generates internal skeletons for existing 3D models, serving as static frameworks with some movement flexibility (Liu et al., 2017). Finally, Bend-it bends kinematic structures by taking key frames of a virtual structure designed by wires as input and generating kinematic mechanisms to animate between these frames, utilizing the springiness and dynamic properties of the wire (Xu et al., 2018). Similar to the aesthetic applications, these fabrication techniques are based on conversion of 3D structures and modification of pre-defined parametric models.
@@ 2.2. Fabrication-Aware Design
Fabrication-aware design is a branch of research in computer graphics that considers the constraints and possibilities of fabrication processes in design environments. This approach ensures that resulting artefacts are manufacturable and functional throughout the design process. Fabrication-aware design has implications in many fields of computational fabrication such as laser cutting, 3D printing, and other fabrication processes.
@@
In the context of laser cutting,kyub is a 3D modeling environment that allows users to design models for laser-cutting in 3D while the software handles the conversion to 2D plates by placing joints for the specific material and machine at hand (Baudisch et al., 2019). FastForce safeguards the structural integrity of such laser cut models (Abdullah et al., 2021) and FoolproofJoint adjusts resulting joints to be optimal for assembly (Park et al., 2022). HingeCore leverages both laser cutting and specific material properties to create models that are aesthetically pleasing and easy to assemble (Abdullah et al., 2022). Fabricaide allows users to preview fabrication process constraints (such as issues in nesting or joint layout) during the design phase (Sethapakdi et al., 2021). FlatFab similarly supports design for fabrication for laser cutting by focusing on a specific assembly technique of intersecting planar sections (McCrae et al., 2014).
@@
In the realm of 3D printing, fabrication-aware design tools have also made significant strides. Bridging the gap introduces a method to efficiently design stable supports using bridges and pillars (Dumas et al., 2014). Stress relief detects structural weaknesses and provides tools for users to adjust their model accordingly (Stava et al., 2012). To increase print efficiency, WirePrint offers a rapid prototyping method by generating wireframe structures from existing 3D models (Mueller et al., 2014), and PackMerger optimizes the segmentation and packing of printed parts to save material and print time (Vanek et al., 2014). Others have brought material specific properties to fabrication-aware design, Wang et al. (2015) reduce fabrication time while maintaining visual quality through adaptive slicing and segmentation techniques. Bickel et al. (2010) use data-driven approaches to create complex, heterogeneous materials with customized deformation properties . Schumacher et al. (2015) design micro-structures to achieve variable elasticity using a single stiff material, expanding supported design applications.
@@
While significant advances have been made in developing fabrication-aware design tools for laser cutting and 3D printing, it is not limited to these processes. In textile design, Plushie allows non-professional users to create plush toys by generating 2D patterns from 3D models (Mori and Igarashi, 2007). For inflatable structures, Skouras et al. (2014) developed a tool for optimizing flat panel patterns for inflatable structures, ensuring accurate design through physics-based simulations. In the project Beyond developable, researchers explore using auxetic materials and conformal geometry to create doubly-curved surfaces from flat pieces (Konaković et al., 2016).
@@
Wire bending remains under explored in terms of fabrication-aware design. Miguel et al. (2016) focus on creating structurally stable wire sculptures rather than serving as a design tool. Wire mesh design explores the broader design potential of wire based materials (Garg et al., 2014). Specialized tools for wire bending could unlock new design possibilities and elevate the complexity and functionality of wire-based structures.
2.3. Computer Aided Design in Mixed Reality
@@
Extended Reality (XR) CAD technologies offer advantages for digital fabrication. Tools like pARam, allow in-situ previewing of designs (Stemasov et al., 2024), Mix&Match enables modeling-free personal fabrication with real-world dimensions (Stemasov et al., 2020), and MixFab, integrates virtual CAD environments and real-world objects to create a mixed-reality environment for personal fabrication, to lower the barrier to entry for broad adoption (Weichel et al., 2014). Additionally, commercial products such as Gravity Sketch555https://www.gravitysketch.com/ demonstrate the practical advantage of XR CAD in design, allowing for quick creation of 3D models, even if they lack the precision of traditional CAD tools. Microsoft’s Maquette666https://store.steampowered.com/app/967490/Microsoft_Maquette/, facilitates spatial prototyping by combining various tools such as surface snapping, hull brush, bounding manipulation, to manipulate 3D content.
@@
XR presents challenges as well. One major issue is the difficulty in accurately obtaining environmental information, and users interacting in mid-air resulting in misalignment between the virtual and physical environments. Even though research has begun exploring ways to align virtual environments with the real world, such as SnapToReality, which proposes a semi-automatic alignment strategy for registering virtual and real spaces, their work relies on scene content and environmental constraints (e.g., edges and surfaces) to anchor virtual content (Nuernberger et al., 2016). Another approach, presented by Fotouhi et al. (2020) requires specific headsets and additional equipment.
@@
Another challenge lies in tracking or gesture detection. Efforts to define hand gestures face limitations, such as high similarity among gestures leading to recognition errors and limited available information (Zhou et al., 2016). Some solutions, like high-quality tracking devices, exist but are often prohibitively expensive and not accessible for everyday home use.
@@
Y-AR leverages the strengths of XR CAD, such as the ease of obtaining real-world dimensions, while mitigating its shortcomings. By combining XR CAD with wire as a material, we leverage the springy nature of materials to design mechanisms that are robust to design, measurement, and fabrication inaccuracies, a concept similar to SpringFit (Roumen et al., 2019) used in mechanical engineering. This approach allows us to create adaptable mechanisms in Y-AR that maintain stability and functionality despite misalignment, addressing limitations related to precise alignment.
@@ 3. Walk-through of Y-AR
Figure 3 shows a walk-through of a user making a mount for a scotch tape extruder on a pegboard. (a) The user puts on their Meta Quest 3 and invokes Y-AR. In Passthrough mode, the user selects the type of connectors. (b) Y-AR places a generic pegboard connector and scales this to the detected gesture, by measuring the distance between the fingers. (c) The user then selects another connector using interface shown in (a) and grabs the scotch tape extruder, to generate a mount which fits the size of the handle. (d) The user links the two connectors to one another to complete the design and fabricate (e) the resulting mechanism with the wire bender.
After design, users export their structure. Export generates a CSV file for the wire bender (shown in Figure 4 as well as assembly instructions for the user in VR. Y-AR aims to make structures out of a single wire, however often this is not feasible, upon which Y-AR exports multiple wires to bend and instructions in VR which guide users through the process of assembling the wires into a 3D structure, as shown in Figure 4b. To rigidly connect wire segments, we use a Crimping Tool Kit777https://www.amazon.com/dp/B0C7V6K4X3 (other similar kits will do too). This allows for secure and precise joining of wires by crimping connectors onto the ends of the wire segments.
@@
We make two core contributions: (1) guidelines for fabrication-aware wire bending and (2) an approach for leveraging material springiness to overcome accuracy issues inherent to mid-air interaction and tracking in XR. We present both contributions as separate sections in the paper as well as integrated in the specific implementation of a first CAD tool for wire bending, we call Y-AR.
@@
(1) We systematically explore the design space of connectors made with wire bending, through this we identify technical guidelines for functional wire bending.
(2) The insight to leverage designed springiness to overcome modeling and tracking inaccuracies serves as a solid basis to implement fabrication-aware CAD tools. We demonstrate this at the example of wire bending in Y-AR. To validate the design of Y-AR, we conduct a usability study.
Limitations include the focus on one specific material used for wire bending which is 1.6 mm red nylon coat wire, while the ideas and methods translate to other materials, it would require re-configuration of some tools in Y-AR. Furthermore, while Y-AR avoids cutting wires (resulting in manual assembly), this algorithm is far from optimal. To the best of our knowledge, most wire benders take csv files containing a list of coordinates as input, however, high-end industrial machines may use more advanced proprietary formats, currently not supported by Y-AR. Y-AR currently supports a series of useful modeling tools, however we realize this set is far from exhaustive. We open-source the code base and make it easy to extend to such formats or other wire types.

### v2

@@ Abstract.
Wire bending is a technique used in manufacturing to mass-produce items such as clips, mounts, and braces. Recent advances in programmable wire bending have made this process increasingly accessible for custom fabrication. However, CNC wire benders are controlled using Computer Aided Manufacturing (CAM) software, without design tools, making custom designs challenging to produce. We present Y-AR, a Computer Aided Design (CAD) interface for 3D wire bending. Y-AR uses mixed reality to let designers create clips, mounts, and braces to physically connect objects to their surrounding environment (such as the mount shown in Figure 1c). The interface incorporates springs as design primitives which (1) apply forces to hold objects, and (2) counter-act dimensional inaccuracies inherently caused by mid-air modeling and measurement errors in AR. Springs are a natural design element when working with metal wire-bending given its specific material properties. We demonstrate workflows to design and fabricate a range of mechanisms in Y-AR as well as structures made using free-hand design tools. We found that combining gesture-based interaction with fabrication-aware design principles allowed novice users to create functional wire connectors, even when using imprecise XR-based input. In our usability evaluation, all 12 participants successfully designed and fabricated a functional bottle holder using Y-AR.
@@ 1. Introduction
Wire bending is a common industrial mass-manufacturing process used to create springs, clips, mounts, and braces, among others. In sculpting, wire bending is used to provide internal support for incomplete structures, for example, as an armature. While most wire bending applications are in high-volume, low-mix mass production, nascent CNC wire bending machines enable on-demand wire bending of high-mix, low volume objects. This mode of production is essential for wire-bending in the medical sector, where it is used to create retainers and braces (Alassiry, 2019), as well as in creative domains such as puppetry, where it is used to produce actuated characters (Xu et al., 2018). To date, most of these applications rely on manual workflows where wires are bent by hand.
@@
New machines such as the DIWire by Pensalabs make automated wire bending accessible for personal fabrication (Baudisch and Mueller, 2017). Hobbyists have embraced it to create connectors, holders, signs, marble tracks, and sculptures. While software exists to operate these machines (e.g., VGPNEXT111https://www.blmgroup.com/en-us/software/vgpnext, Bend-Tech222https://www.bend-tech.com/, and BendPro SIM333https://www.currenttech.com/tube-bending-software/bendpro-sim/), they typically inherit design approaches from industrial manufacturing. These tools provide low-level machine controls rather than giving users tools to design parts that can be manufactured with these machines. This makes it difficult for users without technical expertise to design structures for wire bending.
Besides commercial products, researchers have developed software to convert existing 3D models into bend paths (Huang et al., 2023; Hsiao et al., 2018), enabling hobbyists to create intricate wire structures without having to deal with machine-specific CAM controls. While powerful, these tools are conversion tools, offering users little control and requiring round-trip conversions to edit designs. Moreover, they are agnostic to the constraints of wire bending as a process and the unique material properties of wire structures. Recognizing these challenges, some recent research has proposed more interactive alternatives. WireRoom (Yang et al., 2021), for example, approximates 3D geometry using wire bending, but offers users variations to interactively tune the design to accommodate variations. This provides increased control over the aesthetic conversion of design. However, to date, there are no Computer-Aided Design (CAD) tools specifically designed to support wire bending and take advantage of machine and material properties.
@@
We present Y-AR, a fabrication-aware design (Bermano et al., 2017) tool for 3D wire bending in mixed reality, which enables designers to create structures optimized for CNC wire bending. Y-AR overlays 3D structures onto users’ physical environments in mixed reality, allowing design of objects that fit within the designer’s physical environment (such as the mounts, connectors, and braces shown in Figure 2). We explore the design space of connectors, mounts, and hooks based on wire bending. Y-AR is composed of user-facing features to span most of this design space. Y-AR leverages the elastic properties of wires to create mechanisms that apply force to hold objects in place. Once the design is ready to manufacture, Y-AR lets users export machine instructions for CNC wire benders and generates assembly instructions for users. For fine-tuning, users can still adjust the results in the machine-specific CAM software. Besides guided design through our derived primitives, Y-AR implements basic free-form tools (where users draw structures in mid-air) by modeling a custom marble track.
We found that combining gesture-based interaction with fabrication-aware design principles allowed novice users to create functional wire connectors, even when using coarse XR-based input. We conducted a usability evaluation (n=12) of Y-AR, in which participants designed and fabricated a functional bottle holder using Y-AR.
@@ 2. RELATED WORK
This work builds on prior research on wire bending, fabrication-aware design, and computer-aided design in extended reality.
@@ 2.1. Wire Bending
Research on wire bending has focused on both aesthetic and functional applications. For aesthetic approaches, WrapIt created jigs to speed up manual wire bending for jewelry based on 2D patterns (Iarussi et al., 2015). Fabricable 3D Wire Art similarly generates jigs, however, they are based on 3D models and images, enabling more complex designs (Tojo et al., 2024). With the advent of CNC wire bending machines, there has been an increased focus on advanced wire bending, which involves converting 2D line drawings (Hsiao et al., 2018) or 3D models (Huang et al., 2023) to machine instructions for wire bending. Yeh et al. (2022) propose an algorithm to create 3D wire art with the minimum number of continuous wires. In terms of user interfaces for generating 3D wire art, WireRoom offers an interface that provides multiple options for generating wire bending paths (Yang et al., 2021), allowing users to select their preferred result. PM4WireArt offers users more freedom through a parametric modeling interface designed for conceptual wire art design (Li, 2024). They treat wire as rigid rods assembled into structures, resembling welded frameworks. This approach facilitates the creation of complex designs, highlighting a different use of wire in structural art.
Wire bending has also been studied for its structural and mechanical potential in functional applications. For instance, Bend-forming extends the application of wire bending to aerospace engineering (Bhundiya and Cordero, 2023). Their work involves generating truss structures and exploring the concept of deploying wire benders in space to construct large-scale, static structures.
More down-to-earth, WireFab generates internal skeletons for existing 3D models, serving as static frameworks with some movement flexibility (Liu et al., 2017).
Finally, Bend-it bends kinematic structures by taking key frames of a virtual structure designed by wires as input and generating kinematic mechanisms to animate between these frames, utilizing the springiness and dynamic material properties of the wire (Xu et al., 2018).
These fabrication techniques are based on the conversion of 3D structures and the modification of pre-defined parametric models. This enables users to make much more interesting wire-bend content; however, it lacks the ability to really take advantage of the material and machine properties offered by wire bending. In this paper, we argue that a dedicated CAD environment with primitives centered on the relevant machine and material properties will facilitate more accessible workflows and create fabricated results that leverage the springiness and tolerance of wire structures to remain functional despite coarse input or fabrication inaccuracies.
@@ 2.2. Fabrication-Aware Design
Fabrication-aware design is a branch of research in computer graphics that considers the constraints and possibilities of fabrication processes in design environments. This approach helps ensure that the resulting artifacts are manufacturable and functional throughout the design process. Fabrication-aware design informs a range of computational fabrication techniques, including laser cutting, 3D printing, and other process-specific workflows.
@@
In the context of laser cutting, kyub is a 3D modeling environment that allows users to design models for laser cutting in 3D, while the software handles the conversion to 2D plates by placing joints for the specific material and machine in use (Baudisch et al., 2019). FastForce safeguards the structural integrity of such laser cut models (Abdullah et al., 2021) and FoolproofJoint adjusts resulting joints to be optimal for assembly (Park et al., 2022). HingeCore leverages both laser cutting and specific material properties to create models that are aesthetically pleasing and easy to assemble (Abdullah et al., 2022). Fabricaide allows users to preview fabrication process constraints (such as issues in nesting or joint layout) during the design phase (Sethapakdi et al., 2021). FlatFab similarly supports fabrication-aware design for laser cutting by focusing on a specific assembly technique of intersecting planar sections (McCrae et al., 2014).
@@
In the realm of 3D printing, fabrication-aware design tools have also made significant strides. Bridging the gap introduces a method to efficiently design stable supports using bridges and pillars (Dumas et al., 2014). Stress relief detects structural weaknesses and provides users with tools to adjust their model accordingly (Stava et al., 2012). To increase print efficiency, WirePrint offers a rapid prototyping method by generating wireframe structures from existing 3D models (Mueller et al., 2014), and PackMerger optimizes the segmentation and packing of printed parts to save material and print time (Vanek et al., 2014). Others have introduced material-specific properties to fabrication-aware design. For instance, Wang et al. (2015) reduces fabrication time while maintaining visual quality through adaptive slicing and segmentation techniques. Bickel et al. (2010) employ data-driven approaches to develop complex, heterogeneous materials with customized deformation properties. Schumacher et al. (2015) design micro-structures to achieve variable elasticity using a single stiff material, expanding supported design applications.
@@
There has also been progress in fabrication-aware design for textile design and other fabrication methods. In textile design, for example, Plushie allows non-professional users to create plush toys by generating 2D patterns from 3D models (Mori and Igarashi, 2007). For inflatable structures, Skouras et al. (2014) developed a tool that optimizes flat panel patterns, ensuring accurate design through physics-based simulations. In the project Beyond developable, researchers explore the use of auxetic materials and conformal geometry to create doubly-curved surfaces from flat pieces (Konaković et al., 2016).
@@
Wire bending remains underexplored in terms of fabrication-aware design. Miguel et al. (2016) focus on creating structurally stable wire sculptures rather than serving as a design tool. Wire mesh design explores the broader design potential of wire-based materials (Garg et al., 2014). Specialized tools for wire bending could expand the design space to include both aesthetic and functional artifacts, supporting connectors, mounts, and mechanisms that are tailored to the wire’s unique material and machine constraints.
2.3. Computer-Aided Design in Extended Reality
@@
Extended Reality (XR) CAD technologies offer several advantages for digital fabrication. Tools like pARam, allow in-situ previewing of designs (Stemasov et al., 2024), Mix&Match enables modeling-free personal fabrication with real-world dimensions (Stemasov et al., 2020), and MixFab integrates virtual CAD environments with physical objects to create a mixed-reality workspace that lowers barriers to adoption (Weichel et al., 2014). Additionally, commercial products such as Gravity Sketch444https://www.gravitysketch.com/ demonstrate the practical advantage of XR CAD in design, allowing for quick creation of 3D models, even if they lack the precision of traditional CAD tools. Microsoft’s Maquette555https://store.steampowered.com/app/967490/Microsoft_Maquette/, facilitates spatial prototyping by combining tools like surface snapping, hull brushing, and bounding manipulation for working with 3D content.
@@
XR also presents challenges. One major issue is the difficulty in accurately obtaining environmental information, as users often interact in mid-air, resulting in misalignment between virtual and physical environments. Even though researcher has begun exploring ways to align virtual environments with the real world, such as SnapToReality, which proposes a semi-automatic alignment strategy for registering virtual and real spaces, their work relies on scene content and environmental constraints (e.g., edges and surfaces) to anchor virtual content (Nuernberger et al., 2016). Another approach, presented by Fotouhi et al. (2020) enabling users to visualize different perspectives while actively adjusting the pose.
@@
Other challenge lies in tracking or gesture detection. Efforts to define hand gestures face limitations, such as high similarity among gestures leading to recognition errors and limited available information (Zhou et al., 2016). Some solutions, such as high-quality tracking devices, exist but are often prohibitively expensive and inaccessible for everyday home use.
@@
Y-AR leverages the strengths of XR CAD, such as the ease of obtaining real-world dimensions, while mitigating its shortcomings. By combining XR CAD with wire as a material, we take advantage of the wire’s springy nature to design mechanisms that tolerate inaccuracies in design, measurement, and fabrication, conceptually similar to how SpringFit (Roumen et al., 2019) overcomes inaccuracies of laser cutting machines. This approach allows us to create adaptable mechanisms in Y-AR that maintain stability and functionality despite misalignment, addressing limitations related to precise alignment.
@@ 3. Walk-through of Y-AR
Figure 3 shows a walk-through of a user creating a mount for a Scotch tape dispenser on a pegboard (Figure 3). The user begins by putting on a Meta Quest 3 headset and invoking the Y-AR system. Raising their palm brings up the AR menu, where they select the Connector Tool. To create a pegboard connector, the user points their fingers at the pegboard, and Y-AR interprets this gesture to size and position a connector appropriately. Next, holding the tape dispenser itself, the user defines a mount that fits the handle’s size. After placing these connectors, the user switches to the Link Tool to connect them into a unified wire structure.
After design, users export their structure using the export button shown in Figure 3b. Export generates machine instructions for the wire bender (in CSV format, readable by most typical machines), as well as assembly instructions for the user in AR(shown in the Figure 3f). Y-AR aims to create structures from a single wire; however, when infeasible Y-AR exports multiple wires and provides instructions in AR, which guide users through the process of assembling the wires into a 3D structure, as shown in Figure 3g. To rigidly connect wire segments, we use a Crimping Tool Kit666https://www.amazon.com/dp/B0C7V6K4X3 (other similar kits will also work). This allows for secure and precise joining of wires by crimping connectors onto the ends of the wire segments.
@@
We articulate three core contributions that collectively advance fabrication-aware design for wire bending:
@@
We systematically explore the design space of connectors made with wire bending, through which we identify technical guidelines for functional wire bending.
We contribute a method for modeling and fabricating flexible bent-wire connectors that leverage material stiffness and elasticity to compensate for coarse mixed reality inputs. We validated this approach through a usability study with 12 participants.
We demonstrate the first fabrication-aware design tool for wire bending, going beyond merely aesthetic considerations or conversion to wire bending from other CAD environments.
Limitations include that in this paper, we demonstrate examples using one specific material used for wire bending (1.6mm aluminum). Although the ideas and methods translate to other materials, reconfiguration of some tools in Y-AR would be required (indicated in the relative paper sections). Y-AR is a demonstration of what a fabrication-aware wire-bending tool could enable. We open-source the code base and invite the community to continue expanding the vocabulary of tools for wire bending.
5. Design Space
We define the design space of wire-bent connectors along two key dimensions: the geometry of the target object (1D (holes), 2D (edges), and 3D (volumes)), and how forces are exchanged between connector and object (object pulling on the connector, the connector fixing the object in place, or the connector supporting the weight of the object).
While artistic wire-bent structures can exist independently, such as in free-form sculptures or spatial frames, their most compelling quality from an engineering and design perspective is their ability to interact with other physical objects. The spring-like behavior of wire enables adaptable attachments that can grip, wrap, or support objects without requiring high-precision modeling or rigid alignment. This makes wire an ideal material for creating fast, robust adaptations in physical contexts. Here, we lay out the design space of wire-bent springs as connectors. Figure 4 illustrates this space.
The rows of the design space correspond to the geometric features of the interface between the object and its connector.
1D: The wire interacts with a single point or hole, such as inserting into a peg or pulling through a hole.
2D: The wire attaches to an edge or surface, like clipping to the edge of a table.
3D: The wire engages with the object’s volume, wrapping around it or forming an enclosure.
The columns of the design space describe how forces are exchanged between the object and the structure:
Object pulling on connector: such as hooks, loops, or closed eyelets.
Connector fixating object in place: such as clamps, paperclip, or toggle lock.
Connector supporting the weight of the object: The object applies force on the wire-bent structure from the bottom, such as cantilever arms and holders.
5.1. Translation of design space to design patterns
As we explore the design space and corresponding connectors, we derive specific design guidelines that form the basis of implementation for Y-AR and also more broadly inform the design of effective CNC wire-bending.
5.1.1. Combining connectors
Connectors, by design, reduce the degrees of freedom between the different objects they connect. The specific way each example in the design space accomplishes this will help negotiate trade-offs between different types of connectors. These approaches are not independent of one another and can be combined at will. For example, when a passive object applies a force to the connector, this may depend on the angle at which the object is held relative to gravity, whereas when the connector applies a force to the object directly, this dependency is much less profound.
We demonstrate this principle through two examples shown in Figure 5 a) the cup holder that supports it from below while also pressing in from the sides takes advantage of two connector principles simultaneously, and b) a phone clamp that applies force from both the x and z directions.
5.1.2. Fabrication Constraints
Wire bending is a unique fabrication technique, while versatile and powerful it also introduces its own set of specific constraints when fabricating. We characterize these and demonstrate how we circumvent limitations while maximizing design freedom.
Avoid planar bends
. As the complexity of wire structures increases, collision avoidance becomes increasingly crucial (e.g., three consecutive 90-degree bends); the wire may collide with the machine or the extruded wire itself, leading to wire-bending failure. By introducing small rotations at the nodes within planar structures, as shown in Figure 6, reduces the risk of collisions in the plane.
Minimum bending angle to overcome plastic deformation.
To plastically deform wires requires overcoming the elastic regime of the material. To overcome this, we identify a minimum required bending angle and over-extend bends where needed. Details of this evaluation are included in the technical evaluation of this paper.
Linking connectors.
Structurally, the best way to connect connectors is with at least three non-planar wires, providing rigidity by constraining all degrees of freedom. However, given the linear nature of wire-bending, having a single start and end-point reduces design complexity, enabling fast fabrication while occasionally requiring additional support in one degree of freedom.
Y-AR encourages designing structures using as few wire segments as possible to reduce fabrication effort and simplify assembly, ideally forming each functional connector from a single wire with two endpoints (Fig. 7 a). These wire ends are joined to other segments or structures using metal sleeves and a crimping tool kit777https://www.amazon.com/dp/B0C7V6K4X3, forming a complete assembly (Fig.7 c).
This simplification requires careful attention: many wire structures consist of connectors with two endpoints at approximately equal relative heights, which simplifies assembly but introduces sensitivity to orientation. When joining two wire endpoints using crimping (Fig.7c), designers must consider not only the position but also the relative height and orientation of the connector with respect to the object being gripped. Even if endpoints are correctly positioned, a misalignment between the connector’s grasping direction and the intended load direction can cause the assembled structure to twist or fail to hold the object as intended. For example, the structure shown in Fig. 7 d, Fig.7e shows a structure where the connector properly holds an object in the correct orientation, while Fig. 7f shows the same structure with a height misalignment, resulting in a rotated or unstable grasp.
In some designs, such as hook-based connectors, only one end actively engages with another structure while the remaining free end must be positioned thoughtfully to maintain overall balance (Fig.7b). An example assembled structure, showing how crimping is used to join wire segments securely (Fig.7d).